OpenStudy (mind2000):

4 is what percent of 5

OpenStudy (anonymous):

Percents are always out of 100. So, set up a proportion: \[\frac{ 4 }{ 5 } = \frac{ x }{ 100 }\] Cross multiply and get: \[400=5x\] Divide by 5 and we have our answer: \[x=400/5=80\]

OpenStudy (binary3i):

do you know how to calculate % ? suppose you get X out of Y then the % of X is (X/Y)*100 \[X*100 \div Y\]

OpenStudy (anonymous):

Say you got a 4 out of 5 on a quiz and you want to know what % you got, that's 4/5*100

OpenStudy (anonymous):

4/5=0.8 0.8*100=80%